OpenStudy (anonymous):

A fuel is purified by passing it through a clay pipe. Each foot of the clay pipe removes a fixed percentage of impurities in the fuel. Let f(x) represent the amount of pollutants, in tons, left in the fuel after it had passed through x feet of the clay pipe:

OpenStudy (anonymous):

\[f(x) = 10(0.8)^{x}\] What does the number 0.8 represent? (A) Every foot of the pipe removes 20% of the pollutant. (B) Every foot of the pipe removes 80% of the pollutant. (C) The original amount of pollutant present in the fuel was 80 tons. (D) The original amount of pollutant present in the fuel was 20 tons.

OpenStudy (anonymous):

Can someone plz help me ASAP? @Data_LG2 @DanJS @One098 @sammixboo @arabpride

OpenStudy (arabpride):

0.8 is that same as 8/10 [eight tenths] or can be expressed by 80% ~So which of your options makes the most sense in matching that^

OpenStudy (anonymous):

B?

OpenStudy (arabpride):

I think so c: -makes the most sense to me Hope im right~

OpenStudy (anonymous):

I was actually thinking that it could have been A. But can you clarify with someone else tho, plz?

OpenStudy (arabpride):

I'm not 100% sure myself tbh, but why does A.) make sense to you? :]

OpenStudy (anonymous):

i dont know but 100% - 80% = 20% meaning 0.2. It just happens to be that I would always have to subtract the number enclosed in parentheses by 100%.

OpenStudy (anonymous):

@pooja195 @PRAETORIAN.10 Can u help or clarify?

pooja195 (pooja195):

After every foot, 0.8 or 80% of the pollutants will remain, so how much was removed?

pooja195 (pooja195):

f(x) = 10(0.8)x

OpenStudy (anonymous):

20%

OpenStudy (anonymous):

So it is B? @pooja195

pooja195 (pooja195):

no its A

OpenStudy (anonymous):

ok thnx

OpenStudy (danjs):

The original amount is found at x = 0, a base raised to the zero power is 1 y(0) = 10(0.8)^0 = 10*1 = 10 The initial amount was 10, that crosses off choices C and D

OpenStudy (danjs):

sorry was gone for a few minutes

OpenStudy (danjs):

every time you increase x by 1, the power on 80/100 increases by 1 at x = 1 , you have 10(8/10) at x=2, you have 10(64/100) x= 3, 10(4096/10000) The fractions are becoming smaller by 20 percent of the last fraction each time

OpenStudy (danjs):

IE y = (1/2)^x for increasing x 1/2 ....1/4.....1/8....1/16.....1/32....1/64......1/2^n for n=1,2,3,4,5.....

OpenStudy (danjs):

remember the base in that other problem that was between 0 and 1, always got closer and closer to zero as you went on down the x axis... same thing in this problem... k im done now :)

OpenStudy (anonymous):

thank you ; )

OpenStudy (danjs):

welcome